OpenStudy (anonymous):

find the equation of the line having the given slope and containing the given point m=8/9(4,-5)

OpenStudy (tad1):

y+5 = 8(x -4)

OpenStudy (anonymous):

Use the Point-Slope formula; \[y-y _{1}=m(x-x _{1})\] Put in the slope for m, and the ordered pair for y1 and x1. Then solve to isolate y by itself, and you have your answer.